I am having issues solving this linear equation with a fraction... Please help with step by step... X=-2
Y=5/2x-5
Y=5/2(-2)-5
Not too sure what the next step is :/

Answer by waynest(281)

y = (5/2)x - 5
y = (5/2)-2 - 5
y = -5 - 5
y = -10